OpenStudy (anonymous):

write the equation in vertex form for a function that has a vertex of (-2, 4) and passes through point (1, -2).

OpenStudy (amistre64):

what is 'vertex' form?

OpenStudy (amistre64):

vertex is just the y = (x)^2 + ...

OpenStudy (amistre64):

we know its off center: y-4 = (x+2)^2 is a good start

OpenStudy (amistre64):

we plug in our point to see if we need to adjust the right hand side

OpenStudy (amistre64):

-2-4 = (1+2)^2 +C -6 = 3^2 +C -6 = 9 + C -15 = C ....................... y-4 = (x+2)^2 -15 or: y = (x+2)^2 -11

OpenStudy (anonymous):

for a given parabolic equation we hav, ax^2+bx+c=y the co-ordinates of the vertex is: (-b/2a,c-b^2/4a)... so now put the given points into this... -b/2a=-2=>b=4a and substituing (1,-2) in the equation we hav, a-2b+c=-2 and also c-4a=4. thus summarising the relations we got, b=4a c-4a=4 -7a+c=-2 solve for a,b nd c ... equations is ready!

OpenStudy (amistre64):

..... needs more \(LaTeX\) :)